OpenStudy (anonymous):

can you please help me on 7=5/6c-8

OpenStudy (anonymous):

Add 8 to both sides, multiply by c, and then divide by 15.\[15=\frac{ 5 }{ 6c }\]\[15c=\frac{ 5 }{ 6 }\]\[c=\frac{ 1 }{ 18 }\]

OpenStudy (anonymous):

how do i check it?

OpenStudy (anonymous):

Substitute 1/18 for c, simplify denominator, and multiply by reciprocal \[\frac{ 5 }{ 6c }\] equals five over one third (6 divided by eighteen). Just flip the fraction on the bottom and you get 15-8=7

OpenStudy (anonymous):

ok thanks can you also help me on this other problem 10=2/7n+4